OpenStudy (anonymous):

What are the square roots of 100? Check all that apply. A. 5 B. 10.5 C. -10 D. 10 E. -5 F. |10|

OpenStudy (goformit100):

D. 10

OpenStudy (anonymous):

thats the only one ?

OpenStudy (anonymous):

D,F

Parth (parthkohli):

\( \color{Black}{\Rightarrow \sqrt{x} = \pm\sqrt{x} }\) \( \color{Black}{\Rightarrow \sqrt{100} = \pm10 = |10|}\)

Parth (parthkohli):

It's C, D, F

OpenStudy (anonymous):

whats the answer ?

Parth (parthkohli):

It's already been told. -10 10 |10|

OpenStudy (anonymous):

how is it -10 & |10| ?

OpenStudy (anonymous):

because \(|10|=10\) and \(10^2=100\)

OpenStudy (anonymous):

also \((-10)^2=100\)